OpenStudy (anonymous):

CAN SOMEONE EXPLAIN TO ME HOW DO THIS ?? My teacher didn't really explain.. _Co(g) + O2(g) --> Co2(g)

OpenStudy (rina.r):

first you have to balance it 2CO + O2 = 2CO2 /2CO(g) / + O2(g) /= CO2(g) so you have a 2:1:1 ratio you know the formula n=m/M moles=mass/molar mass well use this later /2CO(g) / + O2(g) /= CO2(g) mass /3.90g / 3.45g / x we call this x because we are trying to determine the maximum mass of CO2 that can form CO is the limiting reactant so we will use O2 to determine the mass of CO2 back to n=m/M for O2 it is n=3.45g/32g/mol 32 is the Molar mass n=0.1078125 moles of O2 there is a 1:1 ratio of O2:CO2 so there are also 0.1078125 moles of CO2 so using n=m/M for CO2 we can plug in the n and also M (32+12) 0.1078125=m/44 m=4.74grams SO the max amount of CO2 that could form is 4.74 grams. http://answers.yahoo.com/question/index?qid=20110119190039AAFjyS5

OpenStudy (anonymous):

that confused me.. hah